OpenStudy (anonymous):

A diameter of a circle has endpoints P(-10,-2) and Q(4,6). a. Find the center of the circle. b.Find the radius. If your answer is not an integer, express it in radical form. c. Write an equation for the circle.

OpenStudy (anonymous):

First it might be more easy on you if you graph it real quick

OpenStudy (anonymous):

Desmos graphing is an easy solution

OpenStudy (anonymous):

Ideally you would find the distance between the points using the pythagorean theorem

OpenStudy (anonymous):

Yea your right I didnt know what I was talking about

OpenStudy (anonymous):

so 14^2 + 8^2 = 196 + 64 = 260, so the distance between the points is the square root of 260

OpenStudy (anonymous):

now just divide by 2 for the radius

OpenStudy (anonymous):

So the radius would be 130?

OpenStudy (anonymous):

@Epicness

OpenStudy (anonymous):

Plus c says to write an equation for the circle, and idk what to do there /:

OpenStudy (anonymous):

And how do you find the center of the circle